OpenStudy (anonymous):

an equation of a circle is given. a. determine its center. b. determine its radius. c. find two points on the circle. x^2 + y^2 = 94

OpenStudy (anonymous):

So far, I have a. (x,y) b. approx 9.70 or sq rt 94

OpenStudy (experimentx):

center is zero.

OpenStudy (experimentx):

your radius is right.

OpenStudy (experimentx):

put any value of x less than radius and find r

OpenStudy (anonymous):

how did you get 0 for center?

OpenStudy (experimentx):

compare with equation standard equation of circle. http://answers.yahoo.com/question/index?qid=20080411074756AAPqBAO

OpenStudy (anonymous):

oohh ok

OpenStudy (anonymous):

so you plug in a number less than the radius for x? and solve for y?

myininaya (myininaya):

\[(x-0)^2=x^2 ; (y-0)^2=y^2 \] So your equation can be written like this: \[(x-0)^2+(y-0)^2=94\]

OpenStudy (experimentx):

yes, it must be less than radius from the center of circle.

myininaya (myininaya):

\[x,y \in [-\sqrt{94}, \sqrt{94}]\]

OpenStudy (anonymous):

I did x=3 and ended up with (3,0) but it's saying 9..

myininaya (myininaya):

\[x^2+y^2=94\] If x=3 the we have \[3^2+y^2=94\] \[9+y^2=94\] \[y^2=94-9\] \[y^2=85\] \[y=\pm \sqrt{85}\] \[(3, \sqrt{85}) \text{ and } (3, - \sqrt{85}) \text{ are two points on the circle}\]

OpenStudy (anonymous):

oh I was just confused when you put that (x-0)^2 + (y-0)^2 = 94, so I was using that..